OpenStudy (littlebird):

lim as x approaches 2 of (sqrt(27-x)-5)/(sqrt(3-x)-1)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you get \[\frac{0}{0}\] when you plug in 2 for \(x\)?

OpenStudy (littlebird):

Yes

OpenStudy (misty1212):

yeah i guess you do too bad

OpenStudy (misty1212):

you have to rationalize BOTH

OpenStudy (misty1212):

i.e. multiply top and bottom by \[(\sqrt{27-x}+5)(\sqrt{3-x}+1)\]

OpenStudy (misty1212):

leave everything that does not get rationalized in factored form, don't multiply out

OpenStudy (misty1212):

clear what i mean, or no?

OpenStudy (littlebird):

yes i understand

OpenStudy (misty1212):

ok the stuff that rationalizes should cancel nicely then plug in 1 remember, do not multiply the other stuff out just leave it

OpenStudy (littlebird):

wow I just stared at it and realized what you were talking about. Thanks! :)